1. determine the molar mass of aluminum perchlorate in units of grams per mole.

2. what is the mass (in grams) of 3.25 moles of elemental nitrogen?

3. how many molecules are found in 18.0 g of carbon dioxide?

4. A 170.00 g sample of an unknown compound want to UC Davis for elemental analysis. The results stated the unknown common consist of 29.84% sodium, 67.49% chromium, and 72.67% oxygen. Find the Paracle formula of this identified compound.
Answer #1

1. determine the molar mass of aluminum perchlorate in units of grams per mole.

solution:

aluminum perchlorate= AlCl3O12

molar mass per mole = (mass of Al) +3(mass of Cl) + 12(mass of O)

= (26.98) + 3(35.45) + 12(15.99)

= 325.21 g/mol

molar mass of aluminum perchlorate is 325.21 g/mol.

2. what is the mass (in grams) of 3.25 moles of elemental nitrogen?

mass of elemental nitrogen (N) = 14 gram/mol

1 mole = 14 gram

3.25 mole = 14 × 3.25 gram

= 45.5 gram

mass of 3.25 mole of elemental nitrogen is 45.5 gram

3. how many molecules are found in 18.0 g of carbon dioxide?

moler mass of CO2 (carbon dioxide) = 44.01 gram/mol

number of moles = given mass/molar mass

= 18(g) / 44.01(g/mol)

=0.409 moles.

1 mole = 6.022 × 1023 molecules

0.409 mole = 0.409 × 6.022 × 1023 molecules

= 2.46 × 1023 molecules of CO2 in 18 g of Carbon dioxide

4.

A 170.00 g sample of an unknown compound want to UC Davis for elemental analysis. The results stated the unknown common consist of 29.84% sodium, 67.49% chromium, and 72.67% oxygen. Find the Paracle formula of this identified compound.

using given percentage - we can correlate in.grams .

29.84 grams of Na

67.49 grams of Cr

72.67 grams of O

number of moles = mass / molar mass.

for Na = 29.84(g) /23(g/mol)

=1.297 moles

for Cr = 67.49(g)/52(g/mol)

=1.297 moles

for O - 72.67(g)/ 15.99(g/mol)

=4.54 mol

to determine molar ratio-

molar Na= 1.297/1.297 =1

molar of Cr = 1.297/1.297 = 1

molar of O = 4.54/1.297 =3.5

so, Na:Cr:O = 1:1:3.5

multiply each by 2

=2:2:7

so empirical formula is.

Na2Cr2O7